OpenStudy (anonymous):

3z^-4*4*w^4*2y^8*y^-7w^-9*z

OpenStudy (anonymous):

Input: 3 x 4 w^4 2 y^8 z/ w^9 z^4 y^7 Result: 24 y / w^5 z^3 Root: w z = ( < That's a cross = ) 0, y = 0 If you want the "Limit" and "Series representations" or you didn't understand what i didn't wrote, Tale a look at this site: http://www.wolframalpha.com/bing/?i=3z%5e-4*4*w%5e4*2y%5e8*y%5e-7w%5e-9*z&expl=3

OpenStudy (anonymous):

\[3\times 4\times 2\times w^{4-9} y^{8-1}z^{-4+1}\]

OpenStudy (anonymous):

be careful about cutting and pasting in to worlfram. it cannot read your mind and in fact i believe the answer above is wrong

OpenStudy (anonymous):

oh no, it is right

OpenStudy (anonymous):

she is right ...

OpenStudy (anonymous):

yeah i made a mistake , it should be \[3\times 4\times 2\times w^{4-9} y^{8-7}z^{-4+1}\]

OpenStudy (anonymous):

\[24w^{-5}yz^{-3}\] etc

OpenStudy (anonymous):

@Hershey_Kisses & @satellite73 : So now Steve got his answer

OpenStudy (anonymous):

thanks :-)